OpenStudy (anonymous):

(x^2)(x^2y^3) forgot how to do these? please explain?

OpenStudy (anonymous):

just multiply. Your x with your x's

OpenStudy (anonymous):

so x*x=x^2

OpenStudy (anonymous):

x*x*x=x^3

OpenStudy (anonymous):

x^5*x=x^6, here we are just adding the exponents.,

OpenStudy (anonymous):

x^4*x^3=x^7

OpenStudy (anonymous):

if we have x^3*y^3, we will just leave it at that since x and y are not the same.

OpenStudy (anonymous):

but if we had x^4*y^3*x^2, we can combine the x's to get x^6*y^3